A standing wave has the following wave-function: y(x,t) = 0.2 sin(πx) cos⁡(12πt), where x and y are in meters, and t is in seconds. If the length of the string is L = 2 m and it is fixed at both ends, then the harmonic, n, in which the string is vibrating is:


n = 3


n = 5


n = 4


n = 6


n = 2

Clear selection



1
Expert's answer
2021-05-24T12:35:00-0400

Length of string, "L=2\space m"

Comparing the given equation with standard equation

"y=A\sin (kx)\cos(\omega t)"


"k=\pi"

"k=\dfrac{2\pi}{\lambda}"

"\lambda=\dfrac{2\pi}{\pi}"

"\lambda=2\space m"


"n=\dfrac{2L}{\lambda}=\dfrac{2\times2}{2}"

"n=2"